OpenStudy (anonymous):

(2x+1)/(9x+y) = 24 find the value of x in the terms of y and vice versa .

OpenStudy (anonymous):

\[\frac{2x+1}{9x+y}=24\] \[2x+1=24(9x+y)\] \[2x+1=216x+9y\]

OpenStudy (anonymous):

now you can solve more easily for x or for y

OpenStudy (saifoo.khan):

SATELITEEEEEEEEEEEEEEEEE.

OpenStudy (saifoo.khan):

your bike moves now.

OpenStudy (anonymous):

hmm still a sensei i see. too much time on chat, note enough answering questions!

OpenStudy (anonymous):

check my process any1 : \[(2x+1)(9x+y) = 24\] doing cross multiplication \[2x+1 =24(9x+y)\] \[2x+1 = 216x+24y\] \[2x=216x+24y-1\] \[2x-216x=24y-1\] \[-214x = 24y-1\] \[x=(24y-1)/(-214)\] \[x=(-24y+1)/214\]

OpenStudy (anonymous):

yes i bike moves and you have a bobblehead!

OpenStudy (saifoo.khan):

yea, im a bit tired & not im a mood to study! HAHA, right.

OpenStudy (anonymous):

\[2x+1 =24(9x+y)\] \[2x+1 = 216x+24y\] y u wrote 9y ?

OpenStudy (saifoo.khan):

http://www.youtube.com/watch?v=suRsxpoAc5w

OpenStudy (anonymous):

@kushashwa you have it

OpenStudy (anonymous):